# Pair Production in time-dependent Electric field at Finite times

### v4

@@ IV.5.1 Approximate expression for transverse momentum distribution function
In this subsection, we discuss the mathematical origin of the substructures that appear in the TMS. Using the approximate analytical expression of the single-particle distribution function in momentum space, which is derived in section III.3. We find the approximate expression for the transverse momentum distribution function is given by:
$f(p_{\perp},y)$ $\approx|N^{+}(p_{\perp})|^{2}\Bigl{(}\mathrm{C}_{0}(p_{\perp},y)+(1-y)\mathrm{C}_{1}(p_{\perp},y)+(1-y)^{2}\mathrm{C}_{2}(p_{\perp},y)\Bigr{)},$ (107)
@@
$\mathrm{C}_{1}(p_{\perp},y)$ $\approx\exp{\Biggl{(}\frac{p_{\perp}^{2}}{2}\tau\Bigl{(}\frac{2E_{0}^{2}\tau}{1+E_{0}^{2}\tau^{2}}-\frac{E_{0}+\pi}{\sqrt{1+E_{0}^{2}\tau^{2}}}\Bigr{)}\Biggr{)}}\cos{(\Upsilon)}$
$\frac{2yE_{0}(1+E_{0}\tau(E_{0}\tau+\sqrt{1+E_{0}^{2}\tau^{2}}))}{(1+E_{0}^{2}\tau^{2})^{3/2}}e^{\pi\tau(E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}})},$
@@
From Eq.(LABEL:TMScoef12) and (110), we easily see that $\mathrm{C}_{1}(p_{\perp},y)$ and $\mathrm{C}_{2}(p_{\perp},y)$ shows Gaussian-like profile with the cosine function and its behavior depends on the field parameters. We now analyze the $\mathrm{C}_{2}(p_{\perp},y)$ function that shows the complicated substructure that changes with time. From Eq.(110) first term shows a complicated substructure having two peaks at $p_{\perp}\approx\pm 1.5[m]$ and dip at the $p_{\perp}\approx 0.$ This structure does not change with time but the second term plays crucial role whose depends the shape of profile on the time since $\cos{(\Upsilon)}$ factor and its gives multi-peak substructure behavior. To further analyze the substructure behavior of the spectra, we reconsider the argument $\Upsilon$ as outlined in section IV.2.1. For this purpose, we approximate $\Upsilon$ by including terms up to $p_{\perp}^{2}$ and disregarding higher-order terms.
@@
$+\frac{p_{\perp}^{2}\tau}{2\sqrt{1+E_{0}^{2}\tau^{2}}}\ln{\Biggl{(}\frac{(1-y)}{1+E_{0}^{2}\tau^{2}}\Biggr{)}}$ (112)
@@
$f(p_{\perp},y\rightarrow 1)\approx|N^{+}(p_{\perp})|^{2}\mathrm{C}_{0}(p_{\perp})$ (113)
@@
$f(p_{\perp})\approx 2e^{2\pi\tau(E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}})}\exp{\biggl{(}\frac{-\pi p_{\perp}^{2}\tau}{\sqrt{1+E_{0}^{2}\tau^{2}}}\biggr{)}}.$ (114)
@@ V Conclusion
We have conducted a detailed analysis of the electron-positron pair production from the vacuum under the influence of a time-dependent Sauter pulse electric field. We calculate the single-particle distribution function with the help of the exact analytic solution for the mode function. We develop an analytical theory that is valid for the finite times $t>\tau.$ In particular, we find an analytical expression for the distribution function in the power series of the small parameter $(1-y).$ The interesting dynamical features of the momentum distribution function at finite times are attributed to the function that appears in this expansion. We have further elaborated on this aspect of the finite-time behavior of the longitudinal and transverse momentum spectra.
@@
We analyzed the temporal evolution of the single-particle distribution function. Our investigation reveals that the process of transition from initially virtual particles to real particles occurs in three distinct stages, which crucially depend on the longitudinal and transverse modes. Moreover, we have quantified the initiation of the REPP stage, determined by the momentum value, with higher momenta resulting in narrower oscillations in the transient region. We meticulously examined the LMS and TMS to understand the momentum-dependent behavior during these stages. In the LMS, one interesting feature at the beginning of the REPP stage is that the spectrum exhibits an oscillating structure as imprints of a quantum signature at the finite time where the electric field is nearly zero. The two-peaked structure in which the central Gaussian-peak structure has onset oscillation, and this quantum interference pattern evolves and fades away. Based on this observation, we identified three distinct time scales associated with this behavior. Additionally, we investigated the impact of different electric field strengths ($E_{0}$) on the duration of the interference pattern’s formation and disappearance. In the multiphoton regime, for $\gamma=2.5$, we explored the time evolution of LMS in different stages of pair production, and the spectrum shows the splitting of a smooth uni-modal structure into a multi-modal Gaussian structure near the REPP region, after which it merges into a single peak Gaussian profile at the asymptotic time limit.
@@
Utilizing the approximate expressions for the momentum distribution function, we unveiled that the LMS structure mainly comprises three distinct functional behaviors. The second-order term dominantly governs the early times, while combining the first and zeroth-order terms leads to the central peak structure with onset oscillation. The first-order term contributes to oscillations, owing to the presence of a Gaussian envelope with oscillation, but its amplitude diminishes with time, resulting in a smooth spectrum profile in late time.
@@
Furthermore, we examined the role of vacuum polarization, $u(p_{\parallel},t)$, and its counterpart, $v(p_{\parallel},t)$, by plotting their time evolution. In the REPP region, both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit nearly identical oscillations with the same amplitude, and this oscillation amplitude decays for higher $p_{\parallel}$ values. This observation implies that the qualitative nature of the vacuum polarization function is significantly influenced by the longitudinal momentum. To elucidate this further, we plotted the LMS of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ These plots reveal two distinct patterns emerging at the start of the REPP stage. One pattern exhibits a Gaussian-like structure, while the other displays a deformed oscillating profile within the Gaussian envelope. These oscillations, exhibiting varying amplitudes, were confined to a small window of longitudinal momentum ($-1<p_{\parallel}<1$).Over time, the regular oscillations of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ became balanced, as explicitly observed in the plot of the LMS of the created particles for late time, where the oscillating behavior was absent due to the equilibrium between $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ As we recognize the pivotal roles of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ functions in accelerating (creation of $e^{-}$) and decelerating (annihilation of $e^{-}$) electrons, respectively, the observed oscillations in the $f(p_{\parallel},t)$ at finite times can be attributed to the characteristic behavior imprints of these functions, resulting from the acceleration and deceleration of electrons in momentum representation, which manifests as the observed oscillating structure. Next, we also discuss the influence of transverse momentum on LMS, which diminishes the value of $f(p_{\parallel})$ and smooths out the oscillation for higher $p_{\perp}$ values. Moreover, we emphasized that these oscillations, seen in the late REPP stage as transverse momentum increases and also impact the formation time of pairs.
@@
Finally, we study the dynamics of TMS, uncovering interesting results that have received limited attention from other authors, particularly in the context of the Sauter pulsed electric field. We observe fluctuating substructures that undergo regular changes in the QEPP and transient region depending on momentum value. This behavior of substructures during this stage is also influenced by longitudinal momentum, resulting in distinct substructures in the spectrum at a given time. This finding indicates that the longitudinal and transverse momentum is related to each other during particle-antiparticle formation quantitatively and qualitatively. At the beginning of the REPP region, that substructure disappears. The shape of the spectrum becomes an invariable Gauss-like profile and is not affected by the longitudinal momentum but its peak value still depends on it during the REPP stage. This contrast highlights the dynamic evolution of substructures in earlier stages, contrasting with the stable Gaussian-like structure observed in the REPP stage. A similar intriguing behavior in the transverse momentum distribution has been observed in the works by Krajewska et al. Krajewska and Kamiński (2019) and Bechler et al. Bechler et al. (2023), depending on the specifics of the electric field parameters. These studies highlight a fascinating characteristic of the transverse distribution function peak that occurs at non-zero $p_{\perp}$ value depending on the frequency. This observation prompts questions about the feasibility of pair formation through tunneling theory, particularly considering the significant distribution function values at non-zero transverse momentum for low frequencies. For our case, the spectrum also shows the occurrence of peak at non-zero transverse momentum signifies the creation of pairs through the multiphoton process at finite times. Although studying the LMS and TMS of created quasi-particles at various dynamical stages shows a critical dependence on longitudinal and transverse momentum, it also brings attention to the intriguing physics behind the behavior of the vacuum in this finite-time scenario for the pair production process.
@@ Appendix A Appendixes
A.1 Normalization of Analytic Mode function
The normalization constants $N^{(\pm)}$ are chosen such that
$\psi^{(\pm)}_{\bm{p}}(y)=\phi^{(\pm)}_{\bm{p}}(y),$ $y\rightarrow 0.$ (115)
Accordingly we have to investigate the behaviour of the mode functions $\psi^{(\pm)}_{\bm{p}}(y)$ in this limit
$\psi^{(+)}_{\bm{p}}(y)$ $=N^{(+)}y^{k}\left(1-y\right)^{l}{}_{2}\mathcal{F}_{1}\left(a,b,c;y\right)$
$=N^{(+)}e^{k\ln{(y)}}e^{l\ln{(1-y)}}{}_{2}\mathcal{F}_{1}\left(a,b,c;y\right)$
=$\stackrel{{\scriptstyle y\to 0}}{{=}}N^{(+)}e^{k\ln{(y)}}=N^{(+)}y^{k}$ (116)
as ${}_{2}\mathcal{F}_{1}\left(a,b,c;y\to 0\right)=1.$ A similar calculation can be done for $\psi^{(+)}_{\bm{p}}(y),$ which results in
$\psi^{(-)}_{\bm{p}}(y)$ =$\stackrel{{\scriptstyle y\to 0}}{{=}}N^{(+)}e^{k\ln{(y)}}=N^{(-)}(\bm{p})y^{-k}$ (117)
It is possible to calculate $\phi^{(\pm)}_{\bm{p}}(y)$ in the same limit. From the definition of the functions , we have
$\phi_{\bm{p}}^{(\pm)}(y)=\frac{e^{\mp\textrm{i}\Theta_{\bm{p}}(y)}}{\sqrt{2\omega(\bm{p},y)\bigl{(}\omega(\bm{p},y)\mp P(p_{\parallel},t)\bigr{)}}}$ (118)
with the transformed dynamical phase $\Theta_{\bm{p}}(y).$ Using the $\partial_{t}=\frac{2}{\tau}y(1-y)\partial_{y}$ we find
[…]
which takes the form
[…]
gives the dynamical phases accumulated between the initial and final states.
The integration can be performed analytically
[…]
$+\omega_{0}(\ln{(y)-\ln{(1+p_{\perp}^{2}+(p_{\parallel}-E_{0}\tau)(p_{\parallel}+(2y-1)E_{0}\tau)+\omega_{0}\omega(\bm{p},y))}})$
$-\omega_{1}(\ln{(1-y)}-\ln{(1+p_{\perp}^{2}+(p_{\parallel}+E_{0}\tau)(p_{\parallel}+(2y-1)E_{0}\tau))}+\omega_{1}\omega(\bm{p},y))$ (121)
In the limit $y\rightarrow 0,$ the dynamical phase splits into a regular part $\tilde{\Theta}_{\bm{p}}(y=0)$ as well as into a divergent part
$\Theta_{\bm{p}}(y)\stackrel{{\scriptstyle y\to 0}}{{=}}\frac{\tau}{2}\omega_{0}\ln{y}+\tilde{\Theta}_{\bm{p}}(0)$ (122)
The mode functions $\phi_{\bm{p}}^{(\pm)}(y)$ reads
$\phi_{\bm{p}}^{(\pm)}(y)=\frac{e^{\mp\textrm{i}\tilde{\Theta}_{\bm{p}}(0)}}{\sqrt{2\omega(\bm{p},0)\bigl{(}\omega(\bm{p},0)\mp P(p_{\parallel},0)\bigr{)}}}e^{\mp\textrm{i}\frac{\tau}{2}\omega_{0}\ln{(y)}}$ (123)
Denoting $P_{0}=P(p_{\parallel},0),\omega_{0}=\omega(\bm{p},0).$ Comparison of (116),(117) and (123) and fixes the normalization constants
$N^{(+)}(\bm{p})=\frac{e^{-\textrm{i}\tilde{\Theta}_{\bm{p}}(0)}}{\sqrt{2\omega_{0}\bigl{(}\omega_{0}-P_{0}\bigr{)}}}$ (124)
$N^{(-)}(\bm{p})=\frac{e^{\textrm{i}\tilde{\Theta}_{\bm{p}}(0)}}{\sqrt{2\omega_{0}\bigl{(}\omega_{0}+P_{0}\bigr{)}}}$ (125)
A.2 Evaluation simple analytical expression for $f(\bm{p},t)$
To investigate the behavior of the distribution functions in the limit $t>>\tau,$ we employ approximations based on the suitable expressions of the Gamma and the Gauss-hypergeometric functions. These approximations enable us to deduce a simplified analytical expression for particle distribution functions. First, we start with approximating the Gauss-hypergeometric function as $y\to 1$. It is crucial to ensure smooth convergence towards the limit of ${}_{2}\mathcal{F}_{1}\left(a,b,c;y\rightarrow 1\right)$. This task is complicated by the intricate nature of the parameters $a$, $b$, and $c$ in this specific context, which makes it essential to exercise caution when dealing with this limit. Therefore, it is beneficial to transform the argument by substituting $y$ with $(1-y)$. This transition can be achieved using the following mathematical identity.
${}_{2}\mathcal{F}_{1}\left(a,b,c;z\right)=\frac{\Gamma\left(c\right)\Gamma\left(c-a-b\right)}{\Gamma\left(c-a\right)\Gamma\left(c-b\right)}{}_{2}\mathcal{F}_{1}\left(a,b,a+b-c+1;1-z\right)\\ +\left(1-z\right)^{c-a-b}\frac{\Gamma\left(c\right)\Gamma\left(a+b-c\right)}{\Gamma\left(a\right)\Gamma\left(b\right)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b+1;1-z\right).\\ \left\lvert\textrm{arg}(1-z)\right\rvert<\pi$ (126)

### v5

@@ IV.5.1 Approximate expression for transverse momentum distribution function
In this subsection, we discuss the mathematical origin of the substructures that appear in the spectrum. Using the approximate analytical expression of the distribution function which is derived in section III.3. We find that the approximate expression for the transverse momentum distribution function is given by:
$f(p_{\perp},y)$ $\approx|N^{(+)}(p_{\perp})|^{2}\Bigl{(}\mathrm{C}_{0}(p_{\perp},y)+(1-y)\mathrm{C}_{1}(p_{\perp},y)+(1-y)^{2}\mathrm{C}_{2}(p_{\perp},y)\Bigr{)},$ (127)
@@
$\mathrm{C}_{1}(p_{\perp},y)$ $\approx\exp\left\{\frac{p_{\perp}^{2}}{2}\tau\left(\frac{2E_{0}^{2}\tau}{1+E_{0}^{2}\tau^{2}}-\frac{E_{0}+\pi}{\sqrt{1+E_{0}^{2}\tau^{2}}}\right)+\pi\tau\left(E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}}\right)\right\}\cos(\Upsilon)$
×$\times\frac{2yE_{0}\left(1+E_{0}\tau(E_{0}\tau+\sqrt{1+E_{0}^{2}\tau^{2}})\right)}{(1+E_{0}^{2}\tau^{2})^{3/2}}.$ (130)
@@
From Eq.(130) and (131), we easily see that $\mathrm{C}_{1}(p_{\perp},y)$ and $\mathrm{C}_{2}(p_{\perp},y)$ shows Gaussian-like profile with the cosine function and its behavior depends on the field parameters. We now analyze the $\mathrm{C}_{2}(p_{\perp},y)$ function that shows the complicated substructure that changes with time. From Eq.(131) first term shows a complicated substructure having two peaks at $p_{\perp}\approx\pm 1.5[m]$ and dip at the $p_{\perp}\approx 0.$ This structure does not change with time but the second term plays crucial role whose depends the shape of profile on the time since $\cos{(\Upsilon)}$ factor and its gives multi-peak substructure behavior. To further analyze the substructure behavior of the spectra, we reconsider the argument $\Upsilon$ as outlined in section IV.2.1. For this purpose, we approximate $\Upsilon$ by including terms up to $p_{\perp}^{2}$ and disregarding higher-order terms.
@@
$+\frac{p_{\perp}^{2}\tau}{2\sqrt{1+E_{0}^{2}\tau^{2}}}\ln{\Biggl{(}\frac{(1-y)}{1+E_{0}^{2}\tau^{2}}\Biggr{)}}.$ (133)
@@
$f(p_{\perp},y\rightarrow 1)\approx|N^{(+)}(p_{\perp})|^{2}\mathrm{C}_{0}(p_{\perp})$ (134)
@@
$f(p_{\perp})$ $\approx 2\exp\left\{2\pi\tau\left(E_{0}\tau-\sqrt{1+E_{0}^{2}\tau^{2}}\right)-\pi\tau\frac{p_{\perp}^{2}}{\sqrt{1+E_{0}^{2}\tau^{2}}}\right\}.$ (135)
@@ V Conclusion
We have conducted a detailed analysis of the electron-positron pair production from the vacuum under the influence of a time-dependent Sauter pulse electric field. We calculated the single-particle distribution function using the exact analytic solution for the mode function. We develop an analytical theory that is valid for the finite times $t>\tau.$ In particular, we find an analytical expression for the distribution function in the power series of the small parameter $(1-y).$ The interesting dynamical features of the momentum distribution function at finite times are attributed to the function that appears in this expansion. We have further elaborated on this aspect of the finite-time behavior of the longitudinal and transverse momentum spectra.
@@
We analyzed the temporal evolution of the single-particle distribution function. Our investigation reveals that the process of transition from initially virtual particles to real particles occurs in three distinct stages, which crucially depend on the longitudinal and transverse momentum components. Moreover, we have quantified the initiation of the REPP stage, determined by the momentum value, with higher momenta resulting in narrower oscillations in the transient region. We meticulously examined the LMS and TMS to understand the momentum-dependent behavior during these stages. In the LMS, one interesting feature at the beginning of the REPP stage is that the spectrum exhibits an oscillating structure as imprints of a quantum signature at the finite time when the electric field is nearly zero. The two-peaked structure in which the central Gaussian-peak structure has onset oscillation, and this quantum interference pattern evolves and fades away. Based on this observation, we identified three distinct time scales associated with this behavior. Additionally, we investigated the impact of different electric field strengths ($E_{0}$) on the duration of the interference pattern’s formation and disappearance. In the multiphoton regime, for $\gamma=2.5$, we explored the time evolution of LMS in different stages of pair production, and the spectrum shows the splitting of a smooth uni-modal structure into a multi-modal Gaussian structure near the REPP region, after which it merges into a single peak Gaussian profile at the asymptotic time limit.
@@
Utilizing the approximate expressions for the momentum distribution function, we showed that the LMS structure arises from three distinct functional behaviors. The second-order term dominantly governs the early times, while combining the first and zeroth-order terms leads to the central peak structure with onset oscillation. The first-order term contributes to oscillations, owing to the presence of a Gaussian envelope with oscillation, but its amplitude diminishes with time, resulting in a smooth spectrum profile in late time.
@@
We examined the role of vacuum polarization $u(p_{\parallel},t)$, and its counterpart $v(p_{\parallel},t)$ by plotting their time evolution. In the REPP region, both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit nearly identical oscillations with the same amplitude, and this oscillation amplitude decays for higher $p_{\parallel}$ values. This observation implies that the qualitative nature of the vacuum polarization function is significantly influenced by the longitudinal momentum. To elucidate this further, we plotted the LMS of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ These plots reveal two distinct patterns emerging at the start of the REPP stage. One pattern exhibits a Gaussian-like structure, while the other displays a deformed oscillating profile within the Gaussian envelope. These oscillations, exhibiting varying amplitudes, were confined to a small window of longitudinal momentum ($-1<p_{\parallel}<1$).Over time, the regular oscillations of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ became balanced, as explicitly observed in the plot of the LMS of the created particles for late time, where the oscillating behavior was absent due to the equilibrium between $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ As we recognize the pivotal roles of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ functions in accelerating (creation of $e^{-}$) and decelerating (annihilation of $e^{-}$) electrons, respectively, the observed oscillations in the $f(p_{\parallel},t)$ at finite times can be attributed to the characteristic behavior imprints of these functions, resulting from the acceleration and deceleration of electrons in momentum representation, which manifests as the observed oscillating structure. Next, we also discuss the influence of transverse momentum on LMS, which diminishes the value of $f(p_{\parallel})$ and smooths out the oscillation for higher $p_{\perp}$ values. Moreover, we emphasized that these oscillations, seen in the late REPP stage as transverse momentum increases and also impact the formation time of pairs.
@@
Finally, we study the dynamics of TMS, uncovering interesting results that have received limited attention from other authors, particularly in the context of the Sauter pulsed electric field. We observe fluctuating substructures that undergo regular changes in the QEPP and transient region depending on the momentum value. This behavior of substructures during this stage is also influenced by longitudinal momentum, resulting in distinct substructures in the spectrum at a given time. This finding indicates that the longitudinal and transverse momentum is related to each other during particle-antiparticle formation quantitatively and qualitatively. At the beginning of the REPP region, that substructure disappears. The shape of the spectrum becomes an invariable Gauss-like profile and is not affected by the longitudinal momentum, but its peak value still depends on it during the REPP stage. This contrast highlights the dynamic evolution of substructures in earlier stages, contrasting with the smooth Gaussian-like profile observed in the REPP stage. A similar intriguing behavior in the transverse momentum distribution has been observed in the works by Krajewska et al. Krajewska and Kamiński (2019) and Bechler et al. Bechler et al. (2023), depending on the specifics of the electric field parameters. These studies highlight a fascinating characteristic of the transverse distribution function peak that occurs at non-zero $p_{\perp}$ value depending on the frequency. This observation prompts questions about the feasibility of pair formation through tunneling theory, particularly considering the significant distribution function values at non-zero transverse momentum for low frequencies. For our case, the spectrum also shows the occurrence of a peak at non-zero transverse momentum signifies the creation of pairs through the multiphoton process at finite times.
Although studying the LMS and TMS of created quasi-particles at various dynamical stages shows a critical dependence on longitudinal and transverse momentum, it also brings attention to the intriguing physics behind the behavior of the vacuum in this finite-time scenario for the pair production process.
@@ Appendix A Appendixes
A.1 Evaluation approximate analytical expression for the distribution function
We have obtained an analytical expression for the distribution function.
$f(\bm{p},y)$ $=|N^{(+)}(\bm{p})|^{2}\Biggl{(}\frac{\omega(\bm{p},y)+P(p_{\parallel},y)}{\omega(\bm{p},y)}\Biggr{)}\bigg{|}\frac{2}{\tau}y(1-y)\frac{ab}{c}f_{1}+\textrm{i}\biggl{(}\omega(\bm{p},y)-(1-y)\omega_{0}-y\omega_{1}\biggr{)}f_{2}\bigg{|}^{2},$ (136)
where
$f_{1}={}_{2}\mathcal{F}_{1}\left(1+a,1+b,1+c;y\right),$ (137)
$f_{2}={}_{2}\mathcal{F}_{1}\left(a,b,c;y\right).$ (138)
We aim to derive an approximate expression to study the behavior of the distribution functions in the limit $t\gg\tau$ (i.e., $y\rightarrow 1$). This is achieved by applying appropriate asymptotic approximations for the Gamma and Gauss-hypergeometric functions, which lead to a simplified analytical form of the particle distribution functions.
To evaluate the limit $y\to 1$ of the hypergeometric function ${}_{2}F_{1}(a,b,c;z)$, we apply appropriate linear transformation formulas Abramowitz and Stegun (1964)
${}_{2}\mathcal{F}_{1}\left(a,b,c;z\right)=\frac{\Gamma\left(c\right)\Gamma\left(c-a-b\right)}{\Gamma\left(c-a\right)\Gamma\left(c-b\right)}{}_{2}\mathcal{F}_{1}\left(a,b,a+b-c+1;1-z\right)\\ +\left(1-z\right)^{c-a-b}\frac{\Gamma\left(c\right)\Gamma\left(a+b-c\right)}{\Gamma\left(a\right)\Gamma\left(b\right)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b+1;1-z\right),\\ \left\lvert\textrm{arg}(1-z)\right\rvert<\pi$ (139)
Using the above relation, we approximate the Gauss-hypergeometric functions $f_{1}$ present in the relation of the distribution function as follows:
$f_{1}$ $={}_{2}\mathcal{F}_{1}\left(1+a,1+b;1+c;y\right)$
$+\frac{\Gamma(1+c)\Gamma(1+a+b-c)}{\Gamma(1+a)\Gamma(1+b)}(1-y)^{(c-a-b-1)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b;1-y\right)$ (140)